a number is 5 less than three times another number. the sum of two numbers is 31. Wat r d numbers?
Answer by Maths68(1220)

Let
x = larger number
y = smaller number

Given
The sum of two numbers is 31
x+y=31
x=31-y..........(1)
A number is 5 less than three times another number
x=3y-5..........(2)
Put the value of x from (2) to (1)
3y-5=31-y
3y+y=31+5
4y=36
4y/y=36/4
y=9
Put the value of y in (1)
x=31-y
x=31-9
x=22
Numbers are 22 and 9